The real rate of interest is 3 percent and inflation is 4 percent. What is the nominal rate of interest?

Multiple Choice

A. 3.00 percent

B. 1.00 percent

C. 7.12 percent

D. -1.00 percent
Answer #1

(1+real rate)=(1+nominal rate)/(1+inflation rate)

(1+0.03)=(1+nominal rate)/(1+0.04)

1+nominal rate=(1.03*1.04)

nominal rate=(1.03*1.04)-1

=7.12%(Approx).